# Quantum Mycielskians: symmetries, twin vertices and distinguishing labelings

### v1

@@ 3.2. Quantum twin vertices in small classical graphs
There is no graph $G$ with $|V(G)|\leq 5$ such that $G$ has quantum symmetries, has no twins, and $\det A_{G}=0$.
@@
$A_{G_{1}^{5}}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{G_{2}^{5}}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&1&1\\ 0&0&1&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{G_{3}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&0&1&1\\ 0&0&0&1&1\\ 1&1&1&0&0\\ 1&1&1&0&0\\ \end{bmatrix},$
$A_{G_{4}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&0&1&1\\ 0&0&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{G_{5}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&0&1\\ 0&1&0&0&1\\ 1&0&0&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{G_{6}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&1&1\\ 0&1&0&1&1\\ 1&1&1&0&0\\ 1&1&1&0&0\end{bmatrix},$
$A_{G^{5}_{7}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&1&1\\ 0&1&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{G^{5}_{8}}=\begin{bmatrix}0&0&1&1&1\\ 0&0&1&1&1\\ 1&1&0&0&1\\ 1&1&0&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{G^{5}_{9}}=\begin{bmatrix}0&0&1&1&1\\ 0&0&1&1&1\\ 1&1&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\end{bmatrix}.$
Among them, only $K_{5}$, $G_{5}^{5}$ and $G_{7}^{5}$ have no twins. However, we have $\det A_{K_{5}}=4$, $\det A_{G_{5}^{5}}=-4$ and $\det A_{G_{7}^{5}}=-2$. ∎
@@
For graphs with $|V(G)|\leq 6$, the only one (up to classical isomorphism) that has both quantum symmetries, no twins, and such that $\det A_{G}=0$ is described by the following adjacency matrix
$A_{G_{6}}=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&1\\ 1&0&0&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}.$
@@
First, by the Fulton criterion ([12], see also [15, Lemma 2.6]) and the fact that for all $1\leq i\leq 6$ we have $\sum\limits_{k=1}^{6}u_{ik}=\sum\limits_{k=1}^{6}u_{ki}=1$, we know that the generating matrix $u$ of $\operatorname{C}(\operatorname{QAut}(G_{6}))$ is of the form
@@
Suppose there exists $\rho_{1},\rho_{2}$, a pair of different actions of a certain quantum group $\mathbb{H}$ on $G_{6}$ such that $\rho_{1}\circ A_{G_{6}}=\rho_{2}\circ A_{G_{6}}$. Let $P$ and $Q$ be matrix representations of $\rho_{1}$ and $\rho_{2}$ on the standard basis of $L^{2}(G)$, respectively. They can be parameterized as
@@
respectively, where $p_{1},p_{2},p_{3}$ and $q_{1},q_{2},q_{3}$ are orthogonal projections in $\operatorname{C}(\mathbb{H})$. Since they must commute with the adjacency matrix, we get $p_{3}=p_{1}$ and $q_{3}=q_{1}$. With this condition, the relation $\rho_{1}\circ A_{G_{6}}=\rho_{2}\circ A_{G_{6}}$ leads to $p_{1}=q_{1}$ and $p_{2}=q_{2}$, so $\rho_{1}$ is identical to $\rho_{2}$, a contradiction. ∎
@@
We analyze the properties of such graphs with non-classical quantum symmetries, listed in [17, Tables 1-2]. We first observe that among them, the only ones with the vanishing determinant of the adjacency matrix are $K_{2}\Box K_{5}$, $C_{10}(4)$, $C_{2}\Box C_{4}$ and $C_{12}(5)$. However, $C_{10}(4)$ possesses five pairs of twin vertices, while $C_{12}(5)$ has six such pairs. For the remaining two graphs, we have
$A_{K_{2}\Box K_{5}}=A_{K_{2}}\otimes\mathds{1}_{5}+\mathds{1}_{2}\otimes A_{K_{5}}=\begin{bmatrix}0&1&1&1&1&1&0&0&0&0\\ 1&0&1&1&1&0&1&0&0&0\\ 1&1&0&1&1&0&0&1&0&0\\ 1&1&1&0&1&0&0&0&1&0\\ 1&1&1&1&0&0&0&0&0&1\\ 1&0&0&0&0&0&1&1&1&1\\ 0&1&0&0&0&1&0&1&1&1\\ 0&0&1&0&0&1&1&0&1&1\\ 0&0&0&1&0&1&1&1&0&1\\ 0&0&0&0&1&1&1&1&1&0\end{bmatrix}$
@@
$A_{C_{4}\Box C_{3}}=A_{C_{4}}\otimes\mathds{1}_{3}+\mathds{1}_{4}\otimes A_{C_{3}}=\left[\begin{array}[]{cccccccccccc}0&1&1&1&0&0&0&0&0&1&0&0\\ 1&0&1&0&1&0&0&0&0&0&1&0\\ 1&1&0&0&0&1&0&0&0&0&0&1\\ 1&0&0&0&1&1&1&0&0&0&0&0\\ 0&1&0&1&0&1&0&1&0&0&0&0\\ 0&0&1&1&1&0&0&0&1&0&0&0\\ 0&0&0&1&0&0&0&1&1&1&0&0\\ 0&0&0&0&1&0&1&0&1&0&1&0\\ 0&0&0&0&0&1&1&1&0&0&0&1\\ 1&0&0&0&0&0&1&0&0&0&1&1\\ 0&1&0&0&0&0&0&1&0&1&0&1\\ 0&0&1&0&0&0&0&0&1&1&1&0\\ \end{array}\right].$
@@
We note that the Fulton criterion does not yield any information about the graphs $K_{2}\Box K_{5}$ (the so-called prism having a basis in $K_{5}$, $\mathrm{Pr}(K_{5})$ [3, Definition 4.2]) and $C_{4}\Box C_{3}$ as they are vertex-transitive [15, Remark 2.7]. The other potential candidate is the Tesseract graph with $16$ vertices. Its adjacency matrix is singular, while it does not have classical twins. The classical symmetry group is $H_{4}$, while the quantum symmetry group is $O_{4}^{-1}$, see [4].
Question 3.16.
@@ 4. The distinguishing number
Let $G=(V,E)$ be a finite graph. A $c$-distinguishing labeling of $G$ is a function $\phi\colon V\to\{1,\dotsc,c\}$ such that for every $\sigma\in\operatorname{Aut}(G)\setminus\{\mathrm{id}\}$, there exists $v\in V$ such that $\phi(v)\neq\phi(\sigma(v))$.
The behavior of the distinguishing number and variations of it under the Mycielski transformation was widely studied in recent years; see, e.g. [1, 14]. Here we introduce a version of Definition 4.1 for quantum graphs. To that end, we recall that a quantum graph $\mathcal{G}$ defines an operator system $S_{\scriptscriptstyle\mathcal{G}}\subset\operatorname{B}(L^{2}(\mathcal{G}))$ through $S_{\scriptscriptstyle\mathcal{G}}=\bigl{\{}m(A\otimes{X})m^{*}\,\bigr{|}\bigl{.}\,X\in\operatorname{B}(L^{2}(\mathcal{G}))\bigr{\}}$ (see, e.g. [13, Section 2.4]).
@@
Let $\mathcal{G}$ be an irreflexive quantum graph, $S_{\scriptscriptstyle\mathcal{G}}$ the operator system associated with $\mathcal{G}$ and $\mathscr{H}$ be a Hilbert space. A partition of unity $\{P_{1},\dotsc,P_{c}\}$ in $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ is called a quantum $c$-labeling of $\mathcal{G}$. We say that that a quantum $c$-labeling of $\mathcal{G}$ is a quantum $c$-coloring of $\mathcal{G}$ if
@@
Note that if $\mathcal{G}$ is a classical graph (i.e. the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathcal{G})$ is commutative) and $\mathfrak{P}$ is a classical $c$-labeling (i.e. $\mathscr{H}$ is one dimensional) then it corresponds to labeling of vertices $f\colon{V_{\mathcal{G}}}\to\{1,\dotsc,c\}$ uniquely defined by the condition $P_{i}(v)=1\Leftrightarrow{f(v)}=i$. In this case $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ can be viewed as a quantum group of automorphisms preserving $f$. In particular, the abelianization of $\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}})$ corresponds to the subgroup of $\operatorname{Aut}(\mathcal{G})$ that preserves $f$.
@@
Let $\mathcal{G}$ be a quantum graph. We define the quantum distinguishing number $D(\mathcal{G})$ of $\mathcal{G}$ as the smallest $c$ such that $\mathcal{G}$ admits a distinguishing $c$-labeling.
@@ Acknowledgments
The work of AB was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2111 – 390814868. This research was partially supported by the University of Warsaw Thematic Research Program “Quantum Symmetries”. AB is supported by the Alexander von Humboldt Foundation. PS was supported by the NCN (National Science Center, Poland) grant no. 2022/47/B/ST1/00582.
Furthermore, the authors wish to thank Adam Skalski for the fruitful discussion on the subject of quantum symmetries of quantum graphs.
Appendix: Proof of Proposition 3.12

### v2

@@ 3.2. Quantum twin vertices in small classical graphs
There is no graph $G$ with $|V(G)|\leq 5$ such that $G$ has quantum symmetries, has no twins, and $\det A_{\scriptscriptstyle G}=0$.
@@
$A_{\scriptscriptstyle G_{1}^{5}}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{\scriptscriptstyle G_{2}^{5}}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&1&1\\ 0&0&1&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{\scriptscriptstyle G_{3}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&0&1&1\\ 0&0&0&1&1\\ 1&1&1&0&0\\ 1&1&1&0&0\\ \end{bmatrix}$,
$A_{\scriptscriptstyle G_{4}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&0&1&1\\ 0&0&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\\ \end{bmatrix},\quad A_{\scriptscriptstyle G_{5}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&0&1\\ 0&1&0&0&1\\ 1&0&0&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{\scriptscriptstyle G_{6}^{5}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&1&1\\ 0&1&0&1&1\\ 1&1&1&0&0\\ 1&1&1&0&0\end{bmatrix}$,
$A_{\scriptscriptstyle G^{5}_{7}}=\begin{bmatrix}0&0&0&1&1\\ 0&0&1&1&1\\ 0&1&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{\scriptscriptstyle G^{5}_{8}}=\begin{bmatrix}0&0&1&1&1\\ 0&0&1&1&1\\ 1&1&0&0&1\\ 1&1&0&0&1\\ 1&1&1&1&0\end{bmatrix},\quad A_{\scriptscriptstyle G^{5}_{9}}=\begin{bmatrix}0&0&1&1&1\\ 0&0&1&1&1\\ 1&1&0&1&1\\ 1&1&1&0&1\\ 1&1&1&1&0\end{bmatrix}$.
Among them, only $K_{5}$, $G_{5}^{5}$ and $G_{7}^{5}$ have no twins. However, we have $\det A_{\scriptscriptstyle K_{5}}=4$, $\det A_{\scriptscriptstyle G_{5}^{5}}=-4$ and $\det A_{\scriptscriptstyle G_{7}^{5}}=-2$. ∎
@@
For graphs with $|V(G)|\leq 6$, the only one (up to classical isomorphism) that has both quantum symmetries, no twins, and such that $\det A_{\scriptscriptstyle G}=0$ is described by the following adjacency matrix
$A_{\scriptscriptstyle G_{6}}=\begin{bmatrix}0&0&0&0&1&1\\ 0&0&1&1&0&1\\ 0&1&0&1&0&1\\ 0&1&1&0&1&1\\ 1&0&0&1&0&0\\ 1&1&1&1&0&0\end{bmatrix}$.
@@
For establishing the next result, we need the following lemma.
Lemma 3.13 (Fulton criterion; [14], see also [17, Lemma 2.6]).
For a finite graph $G$, if $\bigl(A_{\scriptscriptstyle G}^{l}\bigr)_{ii}=\bigl(A_{\scriptscriptstyle G}^{l}\bigr)_{jj}$ for some $l\in\mathbb{N}$, then $u_{i,j}$ – the $(i,j)$-entry of the generating matrix of $\operatorname{C}(\operatorname{QAut}(G))$ – vanishes.
@@
First, by the Fulton criterion and the fact that for all $1\leq i\leq 6$ we have $\sum\limits_{k=1}^{6}u_{ik}=\sum\limits_{k=1}^{6}u_{ki}=1$, we know that the generating matrix $u$ of $\operatorname{C}(\operatorname{QAut}(G_{6}))$ is of the form
@@
Suppose there exists $\rho_{1},\rho_{2}$, a pair of different actions of a certain quantum group $\mathbb{H}$ on $G_{6}$ such that $\rho_{1}\circ A_{\scriptscriptstyle G_{6}}=\rho_{2}\circ A_{\scriptscriptstyle G_{6}}$. Let $P$ and $Q$ be matrix representations of $\rho_{1}$ and $\rho_{2}$ on the standard basis of $L^{2}(G)$, respectively. They can be parameterized as
@@
respectively, where $p_{1},p_{2},p_{3}$ and $q_{1},q_{2},q_{3}$ are orthogonal projections in $\operatorname{C}(\mathbb{H})$. Since they must commute with the adjacency matrix, we get $p_{3}=p_{1}$ and $q_{3}=q_{1}$. With this condition, the relation $\rho_{1}\circ A_{\scriptscriptstyle G_{6}}=\rho_{2}\circ A_{\scriptscriptstyle G_{6}}$ leads to $p_{1}=q_{1}$ and $p_{2}=q_{2}$, so $\rho_{1}$ is identical to $\rho_{2}$, a contradiction. ∎
@@
We analyze the properties of such graphs with non-classical quantum symmetries, listed in [20, Tables 1-2]. We first observe that among them, the only ones with the vanishing determinant of the adjacency matrix are $K_{2}\Box K_{5}$, $C_{10}(4)$, $C_{2}\Box C_{4}$ and $C_{12}(5)$. However, $C_{10}(4)$ possesses five pairs of twin vertices, while $C_{12}(5)$ has six such pairs. For the remaining two graphs, we have
$A_{\scriptscriptstyle K_{2}\Box K_{5}}=A_{\scriptscriptstyle K_{2}}\otimes\mathds{1}_{5}+\mathds{1}_{2}\otimes A_{\scriptscriptstyle K_{5}}=\begin{bmatrix}0&1&1&1&1&1&0&0&0&0\\ 1&0&1&1&1&0&1&0&0&0\\ 1&1&0&1&1&0&0&1&0&0\\ 1&1&1&0&1&0&0&0&1&0\\ 1&1&1&1&0&0&0&0&0&1\\ 1&0&0&0&0&0&1&1&1&1\\ 0&1&0&0&0&1&0&1&1&1\\ 0&0&1&0&0&1&1&0&1&1\\ 0&0&0&1&0&1&1&1&0&1\\ 0&0&0&0&1&1&1&1&1&0\end{bmatrix}$
@@
$A_{\scriptscriptstyle C_{4}\Box C_{3}}=A_{\scriptscriptstyle C_{4}}\otimes\mathds{1}_{3}+\mathds{1}_{4}\otimes A_{\scriptscriptstyle C_{3}}=\left[\begin{array}[]{cccccccccccc}0&1&1&1&0&0&0&0&0&1&0&0\\ 1&0&1&0&1&0&0&0&0&0&1&0\\ 1&1&0&0&0&1&0&0&0&0&0&1\\ 1&0&0&0&1&1&1&0&0&0&0&0\\ 0&1&0&1&0&1&0&1&0&0&0&0\\ 0&0&1&1&1&0&0&0&1&0&0&0\\ 0&0&0&1&0&0&0&1&1&1&0&0\\ 0&0&0&0&1&0&1&0&1&0&1&0\\ 0&0&0&0&0&1&1&1&0&0&0&1\\ 1&0&0&0&0&0&1&0&0&0&1&1\\ 0&1&0&0&0&0&0&1&0&1&0&1\\ 0&0&1&0&0&0&0&0&1&1&1&0\\ \end{array}\right].$
@@
We note that the Fulton criterion does not yield any information about the graphs $K_{2}\Box K_{5}$ (the so-called prism having a basis in $K_{5}$, $\mathrm{Pr}(K_{5})$ [4, Definition 4.2]) and $C_{4}\Box C_{3}$ as they are vertex-transitive [17, Remark 2.7]. The other potential candidate is the Tesseract graph with $16$ vertices. Its adjacency matrix is singular, while it does not have classical twins. The classical symmetry group is $H_{4}$, while the quantum symmetry group is $O_{4}^{-1}$, see [5].
Question 3.17.
@@ 4. The distinguishing number
Let $G=(V,E)$ be a finite graph. A $c$-distinguishing labeling of $G$ is a function $\phi\colon V\to\{1,\dotsc,c\}$ such that for every $\sigma\in\operatorname{Aut}(G)\setminus\{\mathrm{id}\}$, there exists $v\in V$ such that $\phi(v)\neq\phi(\sigma(v))$. The distinguishing number $D(G)$ of $G$ is then the smallest $c$ such that $G$ has a labeling that is $c$-distinguishing.
The behavior of the distinguishing number and variations of it under the Mycielski transformation was widely studied in recent years; see, e.g. [2, 16]. Here we introduce a version of Definition 4.1 for quantum graphs. To that end, we recall that a quantum graph $\mathcal{G}$ defines an operator system $S_{\scriptscriptstyle\mathcal{G}}\subset\operatorname{B}(L^{2}(\mathcal{G}))$ through $S_{\scriptscriptstyle\mathcal{G}}=\bigl\{m(A\otimes{X})m^{*}\,\bigr|\bigl.\,X\in\operatorname{B}(L^{2}(\mathcal{G}))\bigr\}$ (see, e.g. [15, Section 2.4]).
@@
Let $\mathcal{G}$ be an irreflexive quantum graph, $S_{\scriptscriptstyle\mathcal{G}}$ the operator system associated with $\mathcal{G}$ and $\mathscr{H}$ be a Hilbert space. A partition of unity $\{P_{1},\dotsc,P_{c}\}$ in $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ is called a quantum $c$-labeling of $\mathcal{G}$. We say that a quantum $c$-labeling of $\mathcal{G}$ is a quantum $c$-coloring of $\mathcal{G}$ if
@@
Note that if $\mathcal{G}$ is a classical graph (i.e. the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathcal{G})$ is commutative) and $\mathfrak{P}$ is a classical $c$-labeling (i.e. $\mathscr{H}$ is one dimensional) then it corresponds to labeling of vertices $f\colon{V_{\mathcal{G}}}\to\{1,\dotsc,c\}$ uniquely defined by the condition $P_{i}(v)=1\Leftrightarrow{f(v)}=i$. In this case $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ can be viewed as the quantum group of automorphisms preserving $f$. In particular, the abelianization of $\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}})$ corresponds to the subgroup of $\operatorname{Aut}(\mathcal{G})$ that preserves $f$.
@@
Let $\mathcal{G}$ be a quantum graph. We define the quantum distinguishing number $D(\mathcal{G})$ of $\mathcal{G}$ as the smallest $c$ such that $\mathcal{G}$ admits a quantum distinguishing $c$-labeling.
@@ Acknowledgments
The work of AB was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2111 – 390814868. This research was partially supported by the University of Warsaw Thematic Research Program “Quantum Symmetries”. AB is supported by the Alexander von Humboldt Foundation. PS was supported by the NCN (National Science Center, Poland) grant no. 2022/47/B/ST1/00582. ICh is supported by the NCN (National Science Center, Poland) grant SONATA-17 no.  UMO-2021/43/D/ST1/01446. Furthermore, the authors wish to thank Adam Skalski for the fruitful discussion on the subject of quantum symmetries of quantum graphs.
5. Appendix: Proof of Proposition 3.12